Restart Night #8 29 Aug 2020 UT


OSA: S. Allanson

ISp: G. Bechetii

Mtn Manager: B. Bingham

ISA: O. Kuhn

Summary:

We did not open all night due to clouds, then some precipitation.

Proposed Plan:

  • LUCI-1 - CRITICAL PRIORITY
  • PFS Alignment Alignment - daytime - into handover if needed - CRITICAL
  • Focus check time permitting (or if weather keeps us closed at sunset)
  • L! Rotator Center (0.5 hr) - high
    • M92, M34, or M52 - fields available for all night
  • L1 Rotator ZP for AGW - (1 hr) - critical
    • Long Slit Method - instructions to be provided by J. Power
      • Suggest doing 3 "masks" to verify the RotZP (based on experience with issues from L2 run)
  • AGW Transform for both LUCIs (use the N30 fieldstop, H or FeII filters) - (2 hr) - high/critical
  • LUCI-1 Imaging check (can run LUCI-2 in parallel to repeat) - high
    • Important to test in and out of field to "stress-test" AGW-1 - does it recover from long throws? and multi-point dithers?
    • Sidereal
      • Dither in-field
        • J1601 (sets around 10pm) - dithering in the field -Kband only
        • NGC 7027 (sets 4am) - cycles through all filters, can stop early
      • Dither out-of-field
        • M52 - (all night) - KJH
        • NGC 40 - (all night) - all filters
    • NS
      • Comet 29/P - (after 10pm) - JHK - need to update for tonight's ephemeris
  • Twilight Flats - low
  • Photometric Standards - (1-2hrs) - medium - only if clear - any seeing

  • AFTER LUCI-1 checkout:

  • LBC-Red
    • Twlight Flats
  • MODS
    • Spectrophotometric stars if clear (and if not clear only if taking data)
    • Photometric Standards - (1 hr) - only if clear - any seeing
    • Mixed Mode LBC-R/MODS-2
    • Skyflats MODS2-Blue
    • Revisit Transient - (1hr) - low-medium priority - not critical for redshift, if initial spectra look interesting this may be useful

Log

UT 02:15 Staying closed at sunset because of clouds. Dave released for LUCI1 for engineering and Jenny checked and fine-tuned the alignment of the FS for the N3.75 camera.

UT 09:10 There have been clouds passing over since sunset and we have not yet opened. Now the clouds are thick and there are droplets on the all-sky.

UT 12:30 Steve put the telescope to idle. I shut down LBCR and put both MODS to sleep.

-- BarryRothberg - 28 Aug 2020
Topic revision: r6 - 29 Aug 2020, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback